• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 18:23:07 on June 9, 2007

    Entry number 204903

    keyword=Beam Time Accounting
    Hall A

    06/09/07 0800-1600
    Experiment #: E03-101
    Shift Leader: G. Ron
    Shift Workers: R. Shenor, N. Bubis

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    0.220.400.000.000.290.000.000.000.91
    Beam Available
    but Not in Use
    0.550.460.950.940.610.290.000.003.81
    Beam Not Available
    or Unacceptable
    0.230.140.050.060.100.711.001.003.29
    Accelerator Con-
    figuration Change
    0.000.000.000.000.000.000.000.000.00
    Experiment
    Ready
    1.000.000.000.000.501.001.001.004.50
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.001.001.001.000.500.000.000.003.50
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.000.000.00
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.001.001.001.000.500.000.000.003.50
    HRSR0.000.000.000.000.000.000.000.000.00
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: LHRS ioc problem at 9am. Problem with HRS hut door controller prevented access. Jack implemented a work-around for the broken door controller. Moeller measurement began 1:30pm. Eugene determined the Wien angle had been set incorrectly. MCC reset the Wien angle, retuned, and the Moeller setup began around 2:40pm.

    Data Entry by: G. Ron/B. Sawatzky